Does sigma (3n^2-n+1)/sqrt(n^7+2n^2+5) converge or diverge using limit comparison test.

Does

sigma (3n^2-n+1)/sqrt(n^7+2n^2+5) converge or diverge using limit comparison test.
Answer #1

0 to e no comparing with 5 an E 31 - nt nan l+212 +5 Here an= 3n2_nt1 choose bna 0312 - am onza an = n2 [ 3 - 1 2 + 1 2 ] 2eus an - nua bo lm (3-1 + 1 he] nao - 3-1 tto Mitte - 3-oto. Ttoto hawa n = 3 = 3 (20, 0), By Limit comparchion series eitheryou have any doubt Dont forget Matting. comment Thankiy me, i will replay. you